### 12 Dec 2023 v2

@@ 4 Differentiating between population and subpopulation goodness
$\operatorname{E}_{\mathrm{S}}(Y|y,\textbf{x},\hat{\theta})=\frac{\sum_{s=1}^{S}N_{s}\mbox{Pr}(Y_{i\in s}=1|y,\textbf{x},\hat{\theta})}{\sum_{s=1}^{S}N_{s}}.$ (20)
@@
  • $X_{1}\mathrm{only}$: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}),$
  • $X_{3}\mathrm{only}$: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{3})}_{j}),$
We can modify (18) to
$\widehat{\operatorname{Error}}_{\mathrm{S,PSIS-LOCO,MRP}}^{2}=\Bigg{(}\frac{1}{\sum_{s=1}^{S}N_{s}}\sum_{s=1}^{S}N_{s}\frac{\sum_{b=1}^{B}w_{s}^{b}\epsilon_{s}^{b}}{\sum_{b=1}^{B}w_{s}^{b}}\Bigg{)}^{2}$ (21)
@@
$\begin{split}\widehat{\operatorname{CRPS}}_{\mathrm{S,MRP}}(y,n,\hat{\theta},\mathbf{x})=\frac{1}{2}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{s=1}^{S}N_{s}g(y_{i\notin s},\textbf{x},\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})}{\sum_{s=1}^{S}N_{s}}\Bigg{\rvert}\Bigg{)}-\\ \sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{s=1}^{S}N_{s}h(y_{i\notin s},\textbf{x},y_{s},n_{s},\hat{\theta}_{RW}^{b})}{\sum_{s=1}^{S}N_{s}}\Bigg{\rvert}\Bigg{)},\end{split}$ (22)
@@
Figure 6 shows a comparison between the mean over levels of the estimated PSIS-LOCO score for each variable and the true mean over levels. Focusing on the two variables that are strongly predictive of the outcome ($X_{2}$ and $X_{4}$, final two columns) we can cleanly distinguish between models with a low mean true score and models with a high mean true score. This was a demonstrated challenge in Kuh et al. (2022) that these new metrics overcome. Of particular interest for the practitioner is that the models with the lowest true score differ for the different variables, which again reinforces the concept proposed by Kuh et al. (2022) that the best model differs based on the overall aim.
Turning our attention to the variables that are less predictive of the outcome ($X_{1}$ and $X_{3}$, first two columns), we see less clear delineation between good and bad models. However, we still retain the respective ordering despite the noisiness of estimates. The respective score of each level could be plotted individually; see supplementary materials.
Figure 6: Comparison the mean of true scores across levels ($y$-axis) against the mean of estimated scores ($x$-axis, alternating vertical facets) for each level in each variable (horizontal and vertical facets). Colour of the point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. These results can be separated into two classes, those estimating levels of X2 and X4 (strongly predictive of outcome) and those estimating levels of X1 and X3 (weakly predictive of outcome).
@@ 5 Reference validation score
To use our scoring decomposition we require at least one observation in each cell. However, in many real-world cases not all cells will be observed in the sample. To overcome this, we follow the approach of Vehtari & Ojanen (2012) and substitute the observed cellwise probability with an estimated probability using a reference model denoted $M_{*}$. In this section we first consider a full reference validation score where we make this substitution for every cell in Section 5.1. This is so we can compare to previous approaches. Then, in Section 5.2 we consider the case where not all cells are observed. In this section we first extend the PSIS LOCO reference validation to this scenario. The we propose a method we call LOCO combined validation, which is a combination of LOCO cross validation and LOCO reference validation. Finally we propose a method to validate the chosen reference model.
@@ 5.1 All cells observed
The reference validation technique follows the approach of Vehtari & Ojanen (2012). We denote the reference model as $M_{*}$, and compare to a candidate model, $M_{c}$, with the overall aim to evaluate if they perform similarly in estimating the population expectation. The reference model can be difficult to propose (as discussed by Vehtari & Ojanen, 2012), but in an MRP context there is often need to choose the smallest/simplest model. In this paper the model with $X_{4}$ could be used almost interchangeably to the model with all predictors. This is important as the demographic variables need to be estimated in the population, which can be imprecise or require additional modelling (e.g., Gellar et al., 2023).
Similarly, although a model with more complex priors might reduce individual error (e.g., Gao et al., 2021), if the target estimand is at the population level, the additional model and computational complexity might not be justified if the estimate is similarly good with a simpler model.
We can apply a reference validation approach by substituting the reference model estimate for the population truth in (6) and (10). In Equation (30) of Vehtari & Ojanen (2012), the expected loss function under the actual belief model $M_{*}$ is decomposed in the variance for a predicted value given $M_{*}$ and the squared difference in expectations between the reference model $M_{*}$ and the proposed model $M_{c}$. As mentioned by Vehtari & Ojanen (2012) the model minimising the squared loss is that which the predicted expectation is closest to that with the reference model. Incorporating this into (6) in an MRP context, we would minimise
$\operatorname{Error}_{\mathrm{MRP}}(M_{c},M_{*})^{2}=(E_{P}(Y|M_{c},y,\textbf{x})-E_{P}(Y|M_{*},y,\textbf{x}))^{2},$ (24)
@@
$\operatorname{Error}_{\mathrm{MRP}}(M_{c},M_{*})^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|M_{c},y,\textbf{x})-\mbox{Pr}(Y_{i\in j}=1|M_{*},y,\textbf{x}))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2}.$ (25)
@@
where G(t) is the CDF for the candidate model and F(t) is the CDF for the reference model and so we use $\mbox{Pr}(Y|M_{*},y,\textbf{x})$ as Y and $\mbox{Pr}(Y|M_{c},y,\textbf{x})$. Subbing these in and simplifying as in previous sections, we get
$\frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{c},y,\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{c}^{\prime},y,\textbf{x}\bigg{)}\Bigg{|}\,+\\ \frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{*},y,\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{*}^{\prime},y,\textbf{x}\bigg{)}\Bigg{|}\,-\\ \sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{c},y,\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{*},y,\textbf{x}\bigg{)}\Bigg{|}\,.$ (27)
Where we have observed cells, it is important to use the LOCO estimate of $\mbox{Pr}(y_{i\in j}=1|M_{c},y,\textbf{x})$, $\mbox{Pr}(y_{i\in j}=1|M_{c},y_{\notin j},\textbf{x})$ to avoid overfitting and more accurately capture the predictive power for unseen observations. Specifically for squared error, we have
$\operatorname{Error}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}\left(Y_{i\in j}=1|M_{c},y_{i\notin j},\textbf{x}\right)-\mbox{Pr}\left(Y_{i\in j}=1|M_{*},y_{i\notin j},\textbf{x}\right))}{N}\Bigg{)}^{2},$ (28)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\operatorname{CRPS}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})=$
$\sum_{b=1}^{B}\Bigg{(}\frac{1}{2B}\ \Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{c},y_{i\notin j},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{c}^{\prime},y_{i\notin j},\textbf{x}\bigg{)}\Bigg{|}\,+$
$\frac{1}{2}\ \Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{*},y_{i\notin j},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{*}^{\prime},y_{i\notin j},\textbf{x}\bigg{)}\Bigg{|}\,-$
$\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{c},y_{i\notin j},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{*},y_{i\notin j},\textbf{x}\bigg{)}\Bigg{|}\,\Bigg{)}.$
The PSIS approximation of the LOCO reference model estimate for each cell can be substituted. For the squared error score, we simply use an importance-weighted estimate for each cell:
$\operatorname{Error}_{\mathrm{MRP},\mathrm{LOCO},\mathrm{PSIS}}(M_{c},M_{*})^{2}=$ (29)
$\!\!\!\!\!\!\!\!\!\!\Bigg{(}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\frac{\sum_{b=1}^{B}w_{M_{c},j}^{b}\mbox{Pr}(Y_{i\in j}=1|M_{c},\hat{\theta}^{b}_{M_{c}}y_{i\notin j},\textbf{x})}{\sum_{b=1}^{B}w_{M_{c},j}^{b}}-\frac{\sum_{b=1}^{B}w_{M_{*},j}^{b}\mbox{Pr}(Y_{i\in j}=1|M_{*},\hat{\theta}^{b}_{M_{*}}y_{i\notin j},\textbf{x})}{\sum_{b=1}^{B}w_{M_{*},j}^{b}}\bigg{)}\Bigg{)}^{2}.$
Figure 7: Comparison of the candidate model LOCO cross validation score (when estimated using Pareto-smoothed importance sampling, $y$-axis) against the LOCO reference validation approach ($x$-axis). Colour of point represents the different candidate models, whilst different points within this represent different simulation iterations. The black line represents an identity line. The strong relationship indicates that when the reference model is well chosen, this closely approximates the PSIS-LOCO score.
For the LOCO reference validation CRPS score we can use the importance-resampled posteriors for reference model $M_{*}$, $\hat{\theta}_{RW,M_{*}}$ and candidate model $M_{c}$, $\hat{\theta}_{RW,M_{c}}$ We then modify as follows: $g_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{c}}^{\prime b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}_{RW,M_{c}}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}_{RW,M_{c}}^{\prime b},\textbf{x})$ and $h_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}_{RW,M_{c}}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}_{RW,M_{*}}^{b},\textbf{x})$.
@@
$\operatorname{CRPS}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})=\frac{1}{2B}\Bigg{(}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}g_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{c}}^{\prime b})\bigg{)}\Bigg{|}\,+$
@@
$\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}g_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{*}}^{b},\hat{\theta}_{RW,M_{*}}^{\prime b})\Bigg{|}\,-\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}h_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})\Bigg{|}\,\Bigg{)}.$
@@
To demonstrate the efficacy of the PSIS LOCO reference validation approach, we let the full model with all covariates be $M_{*}$ and the candidate models be the model without the precision variable, the model without the bias variable, and the model without either. For each we use a PSIS approximation of the predictive error of cell $j$ and apply this to our previous example with all cells observed. Figure 7 demonstrates that the PSIS LOCO reference validation approach and the PSIS LOCO cross validation approach described in Section 3 produce similar estimates.
These results are conditional on choosing an appropriate reference model. The particular example presented is a best case scenario as the reference model aligns closely with the data generating process. In the following section we discuss a technique to validate the reference model and a method for combining cell scores for observed and unobserved cells.

### 12 Apr 2024 v3

@@ 4 Differentiating between population and subpopulation goodness
$\operatorname{E}_{\mathrm{S}}(Y|y,\hat{\theta})=\frac{\sum_{s=1}^{S}N_{s}\mbox{Pr}(Y_{i\in s}|y,\hat{\theta})}{\sum_{s=1}^{S}N_{s}}.$ (15)
@@
  • $X_{1}~{}\mathrm{only}$ : $\mbox{Pr}(y_{i\in j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{i\in j}),$
  • $X_{3}~{}\mathrm{only}$ : $\mbox{Pr}(y_{i\in j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{3})}_{i\in j}),$
We can modify (13) to estimate the subpopulation score with
$\widehat{\operatorname{Error}}_{S}^{(\mathrm{LOCO-CV})}=\frac{1}{\sum_{s=1}^{S}N_{s}}\sum_{s=1}^{S}N_{s}\frac{\sum_{b=1}^{B}w_{s}^{b}\epsilon_{s}^{b}}{\sum_{b=1}^{B}w_{s}^{b}},$ (16)
@@
$\widehat{\operatorname{CRPS}}_{S}^{(\mathrm{LOCO-CV})}(y,n,\hat{\theta}_{RW},\hat{\theta}^{\prime}_{RW})=\\ \frac{1}{BN}\sum_{b=1}^{B}\Bigg{(}\frac{1}{2}\Bigg{\lvert}\sum_{s=1}^{S}N_{s}g_{s}(\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})\Bigg{\rvert}-\Bigg{\lvert}\sum_{s=1}^{S}N_{s}h_{s}(y_{s},n_{s},\hat{\theta}_{RW}^{b})\Bigg{\rvert}\Bigg{)},$ (17)
@@
Figure 3 shows a comparison between the mean over levels of the estimated PSIS-LOCO score for each variable and the true mean over levels. Focusing on the two variables that are strongly predictive of the outcome ($X_{2}$ and $X_{4}$; final two columns) we can cleanly distinguish between models with a low mean true score and models with a high mean true score. This was a demonstrated challenge in Kuh et al. (2023) that these new metrics overcome. Of particular interest for the practitioner is that the models with the lowest true score differ for the different variables, which again reinforces the concept proposed by Kuh et al. (2023) that the best model differs based on the overall aim.
Turning our attention to the variables that are less predictive of the outcome ($X_{1}$ and $X_{3}$; first two columns), we see less clear delineation between good and bad models. However, we still retain the respective ordering despite the noisiness of estimates. The respective score of each level can also be plotted individually; see supplementary materials.
Figure 3: Comparison of the mean of true scores across levels ($y$-axis) against the mean of estimated scores ($x$-axis, alternating vertical facets) for each level in each variable (horizontal and vertical facets). Colour of the point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. These results can be separated into two classes, those estimating levels of X2 and X4 (strongly predictive of outcome) and those estimating levels of X1 and X3 (weakly predictive of outcome).
@@ 5 Reference validation score
To use our scoring decomposition we require at least one observation in each cell. However, in many real-world cases not all cells will be observed in the sample. To overcome this, we follow the approach of Vehtari & Ojanen (2012) and substitute the observed cellwise probability with an estimated probability using a reference model denoted $M_{*}$. To investigate this, we first consider a full reference validation score where we make this substitution for every cell (Section 5.1), which allows us to compare to cross validation. Then we describe a new simulation scenario where not all cells are observed. We first demonstrate PSIS LOCO reference validation, and then propose a LOCO combined validation approach in which we combine cross validation and reference validation. As a consequence of this, we describe a way of validating the chosen reference model.
@@ 5.1 All cells observed
The reference validation technique follows the approach of Vehtari & Ojanen (2012). We denote the reference model as $M_{*}$, and compare to a candidate model, $M_{c}$, with the overall aim to evaluate to rank the $C$ candidate models and identify candidate models that perform similarly to the reference model. The reference model can be difficult to propose (as discussed by Vehtari & Ojanen, 2012), but in an MRP context there is often need to identify if a smaller or simpler model would suffice . In this paper the model with $X_{4}$ could be considered of similar efficacy to a model with all predictors.
We can apply a reference validation approach by substituting the reference model estimate for the population truth in (4) and (5). In Equation (30) of Vehtari & Ojanen (2012), the expected loss function under the actual belief model $M_{*}$ is decomposed in the variance for a predicted value given $M_{*}$ and the squared difference in expectations between the reference model $M_{*}$ and the proposed model $M_{c}$. As mentioned by Vehtari & Ojanen (2012) the model minimising the squared loss is that which the predicted expectation is closest to that with the reference model. Incorporating this into (4) in an MRP context, we would aim to minimise the square of
$\operatorname{Error}^{(\mathrm{REF})}(M_{c},M_{*})=E_{P}(Y|M_{c},\theta_{c})-E_{P}(Y|M_{*},\theta_{*}),$
@@
$\operatorname{Error}^{(\mathrm{REF})}(M_{c},M_{*})=\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}|\theta_{c})-\mbox{Pr}(Y_{i\in j}|\theta_{*}))}{\sum_{j=1}^{J}N_{j}},$ (19)
with $\mbox{Pr}(Y_{i\in j}|\theta_{c})$ shorthand for $\mbox{Pr}(Y_{i\in j}=1|M_{c},\theta_{c},\textbf{x},y)$.
@@
where $F$ and $G$ are the cumulative distribution functions for the reference and candidate models, respectively, and so we use $\mbox{Pr}(Y|\theta_{*})$ as $Y$ and $\mbox{Pr}(Y|\theta_{c})$ as $X$. Plugging these in and simplifying as in previous sections, we get
$\mathrm{CRPS}^{(\mathrm{REF})}(M_{c},M_{*})=\frac{1}{BN}\Bigg{(}\frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(Y_{i\in j}|\theta_{c}^{b})-\mbox{Pr}(Y_{i\in j}|\theta_{c}^{\prime b}\bigg{)}\Bigg{|}\,+$
$\qquad\frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(Y_{i\in j}|\theta_{*}^{b})-\mbox{Pr}(Y_{i\in j}|\theta_{*}^{\prime b}\bigg{)}\Bigg{|}\,-$
$\sum_{b=1}^{B}\Bigg{|}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(Y_{i\in j}|\theta_{c}^{b})-\mbox{Pr}(Y_{i\in j}|\theta_{*}^{\prime b}\bigg{)}\Bigg{|}\,\Bigg{)}.$
We can implement the LOCO estimate of $\mbox{Pr}(Y_{i\in j}|\theta_{c})$, $\mbox{Pr}(y_{i\in j}|\theta_{c},y_{\notin j})$ to avoid overfitting and more accurately capture the predictive power for unseen observations. For squared error, we use the square of
$\begin{split}\operatorname{Error}^{(\mathrm{LOCO-REF})}(M_{c},M_{*})=\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}\left(Y_{i\in j}|M_{c},y_{i\notin j}\right)-\mbox{Pr}\left(Y_{i\in j}|M_{*},y_{i\notin j}\right))}{N},\end{split}$ (20)
@@
$\mathrm{CRPS}^{(\mathrm{LOCO-REF})}(M_{c},M_{*})=\frac{1}{BN}\sum_{b=1}^{B}\Bigg{(}\frac{1}{2}\Bigg{|}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(Y_{i\in j}|\theta_{c}^{b},y_{i\notin j})-\mbox{Pr}(Y_{i\in j}|\theta_{c}^{\prime b},y_{i\notin j}\bigg{)}\Bigg{|}\,+$
$\qquad\frac{1}{2}\Bigg{|}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(Y_{i\in j}|\theta_{*}^{b},y_{i\notin j})-\mbox{Pr}(Y_{i\in j}|\theta_{*}^{\prime b},y_{i\notin j}\bigg{)}\Bigg{|}\,-$
$\Bigg{|}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(Y_{i\in j}|\theta_{c}^{b},y_{i\notin j})-\mbox{Pr}(Y_{i\in j}|\theta_{*}^{\prime b},y_{i\notin j}\bigg{)}\Bigg{|}\,\Bigg{)}.$
The PSIS approximation of the LOCO reference model estimate for each cell can be substituted. For the error score, we simply use an importance-weighted estimate for each cell:
$\widehat{\operatorname{Error}}^{(\mathrm{LOCO-REF})}(M_{c},M_{*})=\\ \frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\frac{\sum_{b=1}^{B}w_{c,j}^{b}\mbox{Pr}(Y_{i\in j}|\hat{\theta}^{b}_{c})}{\sum_{b=1}^{B}w_{c,j}^{b}}-\frac{\sum_{b=1}^{B}w_{*,j}^{b}\mbox{Pr}(Y_{i\in j}|\hat{\theta}^{b}_{*})}{\sum_{b=1}^{B}w_{*,j}^{b}}\bigg{)}.$ (21)
Figure 4: Comparison of the candidate model LOCO cross validation score (when estimated using Pareto smoothed importance sampling, $y$-axis) against the LOCO reference validation approach ($x$-axis). Colour of point represents the different candidate models, whilst different points within this represent different simulation iterations. The black line represents an identity line. The strong relationship indicates that when the reference model is well chosen, this closely approximates the PSIS-LOCO score.
For the LOCO reference validation CRPS score we can use the importance-resampled posteriors for reference model $M_{*}$, $\hat{\theta}_{*,RW}$ and candidate model $M_{c}$, $\hat{\theta}_{c,RW}$. We then modify as follows: $g_{j,\mathrm{REF}}^{\prime}(\hat{\theta}_{c,RW}^{b},\hat{\theta}_{c,RW}^{\prime b})=\mbox{Pr}(Y_{i\in j}|\hat{\theta}_{c,RW})-\mbox{Pr}(Y_{i\in j}|\hat{\theta}_{c,RW}^{\prime b})$ and $h_{j,\mathrm{REF}}^{\prime}(\hat{\theta}_{c,RW}^{b},\hat{\theta}_{*,RW}^{b})=\mbox{Pr}(Y_{i\in j}|\hat{\theta}_{c,RW}^{b})-\mbox{Pr}(Y_{i\in j}|\hat{\theta}_{*,RW}^{b})$.
@@
$\widehat{\operatorname{CRPS}}^{(\mathrm{LOCO-REF})}(M_{c},M_{*})=\frac{1}{BN}\sum_{b=1}^{B}\Bigg{(}\frac{1}{2}\Bigg{|}\sum_{j=1}^{J}N_{j}g^{\prime}_{j,\mathrm{REF}}(\hat{\theta}_{c,RW}^{b},\hat{\theta}_{c,RW}^{\prime b})\Bigg{|}\,+$
$\frac{1}{2}\Bigg{|}\sum_{j=1}^{J}N_{j}g^{\prime}_{j,\mathrm{REF}}(\hat{\theta}_{*,RW}^{b},\hat{\theta}_{*,RW}^{\prime b})\Bigg{|}\,-\Bigg{|}\sum_{j=1}^{J}N_{j}h^{\prime}_{j,\mathrm{REF}}(\hat{\theta}_{c,RW}^{b},\hat{\theta}_{*,RW}^{b})\Bigg{|}\,\Bigg{)}.$ (22)
@@
To demonstrate the efficacy of the PSIS LOCO reference validation approach, we let the full model with all covariates be $M_{*}$ and the candidate models be the model without the precision variable, the model without the bias variable, and the model without either. For each we use a PSIS approximation of the predictive error of cell $j$ and apply this to our previous example with all cells observed. Figure 4 demonstrates that the PSIS LOCO reference validation approach and the PSIS LOCO cross validation approach described in Section 3 produce similar estimates when the appropriate reference model is chosen.